# Gradient Inversion Attack on Graph Neural Networks

### v1

@@ 4 Theoretical Justification
While prior work Geiping et al. (2020) has shown that the input to a fully-connected layer can be recovered from the layer’s gradients, the analysis of GNNs is largely under-explored. Below we will extensively study whether similar conclusions apply to GraphSAGE and GCN and what parts of the graph input can be recovered from the gradients.
@@ 4.1 Node Attacker
Consider a node classification setting, where the inputs to a GNN layer are the target nodal features and the neighboring nodal features denoted as $\mathbf{x}_{v}$ and $\{\mathbf{x}_{v_{j}}\}_{j\in\mathcal{N}_{v}}$. Through theoretical analysis we will show that parts of the input to a GCN or GraphSAGE layer can be recovered from the gradients analytically without solving an iterative optimization problem.
Algorithm 1 GLG (Node Attacker).
1:Input: $2$-layer GNN model $F(\mathbf{x}_{v},\{\mathbf{x}_{v_{j}}\}_{j\in\mathcal{N}_{v}},\mathbf{W})$
2:$\nabla_{\mathbf{W}}\mathcal{L}_{v}$: gradients calculated by training data, $\mathbf{y}_{v}$: True label
3:Output: private data $\mathbf{x}_{v}$
6:for $p\leftarrow 1\textrm{ to }P$ do
7:   ///// /Compute dummy gradients:
8:   $\nabla_{\mathbf{W}}\hat{\mathcal{L}}_{v}=\partial\mathcal{L}(F(\hat{\mathbf{x}}^{p}_{v},\{\hat{\mathbf{x}}^{p}_{v_{j}}\}_{j\in\mathcal{N}_{v}},\mathbf{W}),\mathbf{y}_{v})\partial\mathbf{W}$
9:   Compute $\mathbb{D}^{p}$ as in equation 12
10:   $\hat{\mathbf{x}}_{v}^{p+1}=\hat{\mathbf{x}}_{v}^{p}-\eta\nabla_{\hat{\mathbf{x}}_{v}^{p}}\mathbb{D}^{p},$
11:   for $j\in\mathcal{N}_{v}$ do
13:   end for
14:end for
15:return $\hat{\mathbf{x}}=\hat{\mathbf{x}}_{v}^{P+1}$
@@
For a GraphSAGE defined in equation 1 or a GCN layer defined in equation 3, Node Attacker 1 can recover the aggregated representations of a target node denoted as $\mathbf{x}_{v}^{\text{agg}}$ (see equation 4 with $\mathbf{h}_{\text{agg}}^{0}=\mathbf{x}_{v}^{\text{agg}}$ ) given the gradients of the first layer as $\mathbf{x}_{v}^{\text{agg}}=\nabla_{(\mathbf{W})_{i}}\mathcal{L}_{v}/\nabla_{(\mathbf{b})_{i}}\mathcal{L}_{v}$ for GCN and as $\mathbf{x}_{v}^{\text{agg}}=\nabla_{(\mathbf{W}_{1})_{i}}\mathcal{L}_{v}/\nabla_{(\mathbf{b})_{i}}\mathcal{L}_{v}$ for GraphSAGE provided that $\nabla_{(\mathbf{b})_{i}}\mathcal{L}_{v}\neq 0$.
@@
We show the proof for a GCN layer, the proof for GraphSAGE is similar and is deferred to the appendix. For a GCN layer, the aggregated input at the target node can be recovered from the gradients with respect to the weights of the first layer by writing the following equations
@@
where $(\mathbf{W})_{i}$ denotes the $i$th row of $\mathbf{W}$, $(\mathbf{b})_{i}$ and $(\tilde{\mathbf{h}}_{v})_{i}$ denotes the $i$th element of vectors $\mathbf{b}$ and $\tilde{\mathbf{h}}_{v}$ respectively. It can be concluded from equation 30 that $\mathbf{x}_{v}^{\text{agg}}$ can be recovered as $\nabla_{(\mathbf{W})_{i}}\mathcal{L}_{v}/\nabla_{(\mathbf{b})_{i}}\mathcal{L}_{v}$ as long as $\nabla_{(\mathbf{b})_{i}}\mathcal{L}_{v}\neq 0$. ∎
@@
For a GraphSAGE layer defined in equation 1, Node Attacker 1 can recover the target node features given the gradients of the first layer as $\mathbf{x}_{v}=\nabla_{(\mathbf{W}_{2})_{i}}\mathcal{L}_{v}/\nabla_{(\mathbf{b})_{i}}\mathcal{L}_{v}$ as long as $\nabla_{(\mathbf{b})_{i}}\mathcal{L}_{v}\neq 0$.
@@
In the following, we focus on theoretically analyzing the attack setting of Node Attacker 2, specifically targeting the recovery of nodal features and/or the graph structure. To reiterate, this setting involves leveraging the gradients of the loss function of GNNs associated with each node in the egonet or subgraph.
@@
For a GCN layer defined in equations equation 3 and equation 4, Node Attacker 2-a can recover $\tilde{\mathbf{A}}$ given the gradients of the first layer as long as $\mathbf{X}$ is full row-rank.
@@
From Proposition 1 we know that given the gradients for a node we can recover the aggregated representations of that node. Since we know the gradients for each node we can recover the aggregated representation for each node. Let $\mathbf{X}^{\text{agg}}\in\mathbb{R}^{N\times D}$ denote the aggregated node representation matrix where each row $i$ denotes the aggregated node representation of node $v_{i}$, equation 4 in the matrix form can be written as the following for $l-1=0$.
@@
Since $\mathbf{X}^{\text{agg}}$ and $\mathbf{X}$ are known, the adjacency matrix can be then recovered as $\tilde{\mathbf{A}}=\mathbf{X}^{\text{agg}}\mathbf{X}^{\dagger}$. ∎
@@
For a GCN layer defined in equations equation 4 and equation 3, Node Attacker 2-b can recover $\mathbf{X}$ given the gradients of the first layer if $\tilde{\mathbf{A}}$ is full column-rank.
@@
Algorithm 2 GLG (Node Attacker 2)
2:Output: private training data $\mathbf{X},\mathbf{A}$
4:for $p\leftarrow 1\textrm{ to }P$ do
5:   ///// /Compute dummy gradients:
6:   $\nabla_{\mathbf{W}}\hat{\mathcal{L}}=\partial\mathcal{L}(F(\hat{\mathbf{X}}^{p},\hat{\mathbf{A}}^{p},\mathbf{W}),\mathbf{y})/\partial\mathbf{W}$
7:   Compute $\hat{\mathbb{D}}^{p}$ as given in equation 14
8:   $\hat{\mathbf{X}}^{p+1}=\hat{\mathbf{X}}^{p}-\eta\nabla_{\hat{\mathbf{X}}^{p}}\hat{\mathbb{D}}^{p},$
9:   $\hat{\mathbf{A}}^{p+1}=\text{proj}_{[0,1]}(\hat{\mathbf{A}}^{p}-\eta\nabla_{\hat{\mathbf{A}}^{p}}\hat{\mathbb{D}}^{p}),$
10:end for
@@
For a GraphSAGE layer defined in equation 1, Node Attacker 2-c can recover both $\mathbf{X}$ and $\tilde{\mathbf{A}}$ given only the gradients of the first layer for each node.
@@
Proposition 2 states that for a GraphSAGE layer, the target node features can be recovered given the gradients of the first layer for that node. Similarly, if we have access to the gradients for each node in the graph we can recover all the nodal features. Let $\mathbf{X}$ denote this matrix. From Propositions 1 it is known that for a GraphSAGE layer, the aggregated node representations $\mathbf{X}^{\text{agg}}$ can be recovered. With both $\mathbf{X}$ and $\mathbf{X}^{\text{agg}}$ in hand, the adjacency matrix can be recovered as $\tilde{\mathbf{A}}=\mathbf{X}^{\text{agg}}\mathbf{X}^{\dagger}$ similar to Proposition 3. ∎
The aforementioned proposition reveals big privacy concerns. Even without any prior knowledge of the graph data, the attacker can successfully reconstruct both the nodal features $\mathbf{X}$ and the underlying graph structure $\mathbf{A}$. This result is particularly important as one would anticipate that the intertwined nature of graph data makes it impossible to independently recover $\mathbf{X}$ and $\mathbf{A}$, which is however shown not to be true.
Algorithm 3 Graph Attacker
2:Output: private training data $\mathbf{X},\mathbf{A}$
4:for $p\leftarrow 1\textrm{ to }P$ do
5:   ///// /Compute dummy gradients:
6:   $\nabla_{\mathbf{W}}\hat{\mathcal{L}}=\partial\mathcal{L}(F(\hat{\mathbf{X}}^{p},\hat{\mathbf{A}}^{p},\mathbf{W}),\mathbf{y})/\partial\mathbf{W}$
7:   Compute $D^{p}$ as given in equation 12
8:   $\hat{\mathbf{X}}^{p+1}=\hat{\mathbf{X}}^{p}-\eta\nabla_{\hat{\mathbf{X}}^{p}}\mathbb{D}^{p},$
9:   $\hat{\mathbf{A}}^{p+1}=\text{proj}_{[0,1]}(\hat{\mathbf{A}}^{p}-\eta\nabla_{\hat{\mathbf{A}}^{p}}\mathbb{D}^{p}),$
10:end for
@@ 4.2 Graph Attacker
The present subsection provides theoretical analysis for a Graph Attacker-a in a scenario where prior information may be available for the graph data to be attacked. It can be shown that for GraphSAGE, prior knowledge of the nodal feature matrix $\mathbf{X}$ can help recover the graph structure $\mathbf{A}$, making the framework more vulnerable to the gradient-based attacks.
@@
For a GraphSAGE framework defined in equation 6 with $\mathbf{H}^{0}=\mathbf{X}$, Graph Attacker-a can recover $\tilde{\mathbf{A}}$ from the gradients of the first layer as $\tilde{\mathbf{A}}=\mathbf{X}(\nabla_{\mathbf{W}_{2}}\mathcal{L})^{\dagger}(\nabla_{\mathbf{W}_{1}}\mathcal{L})\mathbf{X}^{\dagger}$ if the nodal feature matrix $\mathbf{X}$ is full-row rank and known.
@@
Therefore if the attacker knows the nodal feature matrix $\mathbf{X}$ along with the gradients, it can recover the normalized adjacency matrix $\tilde{\mathbf{A}}$ analytically.
Even though it is not possible to analytically demonstrate the recovery for Graph Attacker-b and Graph Attacker-c, we do conduct experiments for all the threat models for both GraphSAGE and GCN.
Table 1: Target node feature recovery $\mathbf{x}_{v}$ (RNMSE) for Node Attacker 1.
Synthetic FB GitHub
@@ 5 Data & Experiments
In this section, the performance of the algorithms is tested over real-world social network datasets as well as synthetic datasets to validate the efficacy of the proposed framework.
Evaluation Metrics
Error metrics for recovering the nodal features and the graph structure are evaluated, along with their standard deviation.
  • Nodal Features: To evaluate the performance of the recovered node features, the Root Normalised Mean Squared Error(RNMSE) is used
@@
    In order to evaluate the performance of recovering $\mathbf{X}$ in a graph classification setting, we report the mean RNMSE over all the nodal features in $\mathbf{X}$.
  • Graph Structure: To evaluate the recovery performance of the binary adjacency matrix, we use the following metrics
@@
Algorithms 1, 2 and 3 solve an optimization problem using gradient descent. The gradient descent step can be replaced with off-the-shelf optimizers such as Adam or L-BFGS. In all the experiments, Adam is used as the optimizer. In an attack setting, allowing multiple restarts to the optimizer (especially L-BFGS) from different starting points can significantly increase the quality of recovery. For a fair comparison, multiple restarts are not used, and all results presented are obtained with a single run of the optimizer. However, in an actual attack setting, an attacker may be able to greatly improve the recovery quality by allowing multiple restarts.
In the experiments, a 2-layer GNN model with hidden dimension $100$ and a sigmoid activation function is employed. The weights of the model are randomly initialized. For all the experiments, the dummy nodal features are initialized randomly, with each entry sampled from the standard normal distribution i.e., $\mathcal{N}(0,1)$. The dummy adjacency matrix is initialized by randomly setting its entries to 00 or $1$. The hyperparameters values for the feature smoothness and the sparsity regularizers are set to $\alpha=10^{-9}$ and $\beta=10^{-7}$. These are the values selected by grid search for the choice that leads to the best performance.
@@
All results reported are averaged over $20$ runs of the attacks.
@@
We consider the following three datasets.
  • GitHub Rozemberczki et al. (2021): Nodes represent developers on GitHub and edges are mutual follower relationships.The nodal features are extracted based on the location, repositories starred, employer and e-mail address. Binary labels indicate the user’s job title, either web developer or machine learning developer.
  • FacebookPagePage (FB) Rozemberczki et al. (2019): A page-page graph of verified Facebook sites. Nodes correspond to official Facebook pages, links represent mutual likes between sites. Node features are extracted from the site descriptions. The labels denote the categories of the sites.
@@ 5.1 Node Attacker 1
We first evaluated the performance of feature recovery of Node Attacker 1 for the node classification task.
5.1.1 Target node feature recovery
In this experiment, since no knowledge of the graph data set is available on the attacker side, an initial dummy graph is generated as a 2-layer tree with the target node as the root node and the degree of each node $d_{tree}=10$.
Table 1 lists the RNMSE of recovering the target node features. The results are obtained by running the attack over 20 randomly selected nodes from each data set. It can be observed that for GraphSAGE, the attacker can recover the nodal features with high accuracy. For GCN, the attacker fails to recover the target nodal features. Such results are consistent with the theoretical analysis in Propositions 1 and 2. To further evaluate the performance, we also evaluated the minimum RNMSE obtained in Table 1, which indicates the best recovery performance among the $20$ nodes. It can be observed that while the mean RNMSE of the target node features for GCN is large, the minimum RNMSE can be small. This implies that data might also be leaked when GCN is used in pracitice.
Figure 2: One-hop neighbor recovery (RNMSE) for Node Attacker-1 with GraphSAGE.
Table 2: Recovery of $\mathbf{A}$ with Node Attacker 2-a.
@@ 5.1.1 Target node feature recovery
ACC. AUC AP ACC. AUC AP
GraphSAGE 1.0$\pm$0.0 1.0$\pm$0.0 1.0$\pm$0.0 1.0$\pm$0.0 1.0$\pm$0.0 1.0$\pm$0.0
GCN 0.97$\pm$0.008 0.97$\pm$0.02 0.85$\pm$0.04 0.97$\pm$0.02 0.98$\pm$0.001 0.87$\pm$0.02
Table 3: Recovery of $\mathbf{X}$ with Node Attacker 2-b.
Facebook GitHub
@@
5.1.2 One-hop Neighborhood Nodal Feature Recovery
To evaluate the performance of the node attacker in recovering the nodal features of one-hop neighbors, a target node is selected randomly from the graph. A dummy graph is initialized as a depth-$2$ tree with $d_{tree}$ set to the actual number of one-hop neighbors of the target node. Since there’s permutation ambiguity in the recovered nodal features, the Hungarian Algorithm is used to find the best match with the original neighboring node features Kuhn (2010). Figure 2 shows the one-hop neighbor recovery error (RNMSE) using GraphSAGE. The results are averaged over $10$ randomly selected nodes. For some nodes, the nodal features of neighbors were recovered accurately. For example, nodes 00 and $8$ in Figure 2(a) for Facebook. Further, although the mean RNMSE of the recovered neighbors is high, the minimum RNMSE of the best-recovered neighbors can be low (node $9$ in Figure 2(a)), indicating private data leakage.
Table 4: Recovery of $\mathbf{X}$ (RNMSE) and $\mathbf{A}$ (AUC and AP) with Node Attacker 2-c.
Framework Facebook GitHub
$\mathbf{X}\downarrow$ (RNMSE) AUC $\uparrow$ AP $\uparrow$ $\mathbf{X}\downarrow$ (RNMSE) AUC $\uparrow$ AP $\uparrow$
GraphSAGE 0.002$\pm$0.003 0.99$\pm$0.01 0.99$\pm$0.0004 0.001$\pm$0.002 0.99$\pm$0.004 0.99$\pm$0.009
GCN 1.39$\pm$0.49 0.97$\pm$0.03 0.75$\pm$0.05 0.58$\pm$0.10 0.96$\pm$0.01 0.83$\pm$0.04
GCN (w/o reg.) 1.19$\pm$0.08 0.95$\pm$0.001 0.65$\pm$0.2 0.81$\pm$0.41 0.94$\pm$0.02 0.72$\pm$0.09

### v2

@@ 5 Theoretical Justification
While prior work (Geiping et al., 2020) has shown that the input to a fully-connected layer can be reconstructed from the layer’s gradients, the analysis of GNNs is largely under-explored. Below we will extensively study whether similar conclusions apply to GraphSAGE and GCN and what parts of the graph input can be reconstructed from the gradients.
@@ 5.1 Node Attacker
Consider a node classification setting, where the inputs to a GNN layer are the target nodal features and the neighboring nodal features denoted as $\mathbf{x}_{v}$ and $\{\mathbf{x}_{v_{j}}\}_{j\in\mathcal{N}_{v}}$. Through theoretical analysis we will show that parts of the input to a GCN or GraphSAGE layer can be reconstructed from the gradients analytically without solving an iterative optimization problem.
@@
For a GraphSAGE defined in equation 1 or a GCN layer defined in equation 3, Node Attacker 1 can reconstruct the aggregated representations of a target node denoted as $\mathbf{x}_{v}^{\text{agg}}$ (see equation 4 with $\mathbf{h}_{\text{agg}}^{0}=\mathbf{x}_{v}^{\text{agg}}$ ) given the gradients of the first layer as $\mathbf{x}_{v}^{\text{agg}}=\nabla_{(\mathbf{W})_{i}}\mathcal{L}_{v}/\nabla_{(\mathbf{b})_{i}}\mathcal{L}_{v}$ for GCN and as $\mathbf{x}_{v}^{\text{agg}}=\nabla_{(\mathbf{W}_{1})_{i}}\mathcal{L}_{v}/\nabla_{(\mathbf{b})_{i}}\mathcal{L}_{v}$ for GraphSAGE provided that $\nabla_{(\mathbf{b})_{i}}\mathcal{L}_{v}\neq 0$.
@@
We show the proof for a GCN layer. For a GCN layer, the aggregated input at the target node can be reconstructed from the gradients with respect to the weights of the first layer by writing the following equations
@@
where $(\mathbf{W})_{i}$ denotes the $i$th row of $\mathbf{W}$, $(\mathbf{b})_{i}$ and $(\tilde{\mathbf{h}}_{v})_{i}$ denotes the $i$th element of vectors $\mathbf{b}$ and $\tilde{\mathbf{h}}_{v}$ respectively. It can be concluded from equation 30 that $\mathbf{x}_{v}^{\text{agg}}$ can be reconstructed as $\nabla_{(\mathbf{W})_{i}}\mathcal{L}_{v}/\nabla_{(\mathbf{b})_{i}}\mathcal{L}_{v}$ as long as $\nabla_{(\mathbf{b})_{i}}\mathcal{L}_{v}\neq 0$. ∎
@@
For a GraphSAGE layer defined in equation 1, Node Attacker 1 can reconstruct the target node features given the gradients of the first layer as $\mathbf{x}_{v}=\nabla_{(\mathbf{W}_{2})_{i}}\mathcal{L}_{v}/\nabla_{(\mathbf{b})_{i}}\mathcal{L}_{v}$ as long as $\nabla_{(\mathbf{b})_{i}}\mathcal{L}_{v}\neq 0$.
@@
In the following, we focus on theoretically analyzing the attack setting of Node Attacker 2, specifically targeting the reconstruction of nodal features and/or the graph structure. To reiterate, this setting involves leveraging the gradients of the loss function of GNNs associated with each node in the egonet or subgraph.
@@
For a GCN layer defined in equations equation 3 and equation 4, Node Attacker 2-G can reconstruct $\tilde{\mathbf{A}}$ given the gradients of the first layer as long as $\mathbf{X}$ is full row-rank.
@@
From Proposition 1 we know that given the gradients for a node we can reconstruct the aggregated representations of that node. Since we know the gradients for each node we can reconstruct the aggregated representation for each node. Let $\mathbf{X}^{\text{agg}}\in\mathbb{R}^{N\times D}$ denote the aggregated node representation matrix where each row $i$ denotes the aggregated node representation of node $v_{i}$, equation 4 in the matrix form can be written as the following for $l-1=0$.
@@
Since $\mathbf{X}^{\text{agg}}$ and $\mathbf{X}$ are known, the adjacency matrix can be then reconstructed as $\tilde{\mathbf{A}}=\mathbf{X}^{\text{agg}}\mathbf{X}^{\dagger}$. ∎
@@
For a GCN layer defined in equations equation 4 and equation 3, Node Attacker 2-N can reconstruct $\mathbf{X}$ given the gradients of the first layer if $\tilde{\mathbf{A}}$ is full column-rank.
@@
For a GraphSAGE layer defined in equation 1, Node Attacker 2-GN can reconstruct both $\mathbf{X}$ and $\tilde{\mathbf{A}}$ given only the gradients of the first layer for each node.
@@
Proposition 2 states that for a GraphSAGE layer, the target node features can be reconstructed given the gradients of the first layer for that node. Similarly, if we have access to the gradients for each node in the graph we can reconstruct all the nodal features. Let $\mathbf{X}$ denote this matrix. From Propositions 1 it is known that for a GraphSAGE layer, the aggregated node representations $\mathbf{X}^{\text{agg}}$ can be reconstructed. With both $\mathbf{X}$ and $\mathbf{X}^{\text{agg}}$ in hand, the adjacency matrix can be reconstructed as $\tilde{\mathbf{A}}=\mathbf{X}^{\text{agg}}\mathbf{X}^{\dagger}$ similar to Proposition 3. ∎
The aforementioned proposition has significant privacy implications. Even without any prior knowledge of the graph data, the attacker can successfully reconstruct both the nodal features $\mathbf{X}$ and the underlying graph structure $\mathbf{A}$. This result is particularly important as one would anticipate that the intertwined nature of graph data makes it impossible to independently reconstruct $\mathbf{X}$ and $\mathbf{A}$, which is however shown not to be true.
@@ 5.2 Graph Attacker
The present subsection provides theoretical analysis for a Graph Attacker-G in a scenario where prior information may be available for the graph data to be attacked. It can be shown that for GraphSAGE, prior knowledge of the nodal feature matrix $\mathbf{X}$ can help reconstruct the graph structure $\mathbf{A}$, making the framework more vulnerable to the gradient-inversion attacks.
@@
For a GraphSAGE framework defined in equation 6 with $\mathbf{H}^{0}=\mathbf{X}$, Graph Attacker-G can reconstruct $\tilde{\mathbf{A}}$ from the gradients of the first layer as $\tilde{\mathbf{A}}=\mathbf{X}(\nabla_{\mathbf{W}_{2}}\mathcal{L})^{\dagger}(\nabla_{\mathbf{W}_{1}}\mathcal{L})\mathbf{X}^{\dagger}$ if the nodal feature matrix $\mathbf{X}$ is full-row rank and known.
@@
Therefore if the attacker knows the nodal feature matrix $\mathbf{X}$ along with the gradients, it can reconstruct the normalized adjacency matrix $\tilde{\mathbf{A}}$ analytically.
Even though it is not possible to analytically demonstrate the reconstruction for Graph Attacker-N and Graph Attacker-GN, we do conduct experiments for all the threat models for both GraphSAGE and GCN.
Table 3: Target node feature reconstruction $\mathbf{x}_{v}$ (RNMSE) for Node Attacker 1.
Synthetic FB GitHub OGBN-Arxiv
GLG (Ours)
@@ 6 Data & Experiments
The effectiveness of the proposed algorithms is evaluated on real-world social network datasets, paper citation network datasets, molecular datasets, and synthetic datasets. The detailed results for molecular datasets are provided separately in Appendix G.
Evaluation Metrics.
Error metrics for reconstructing the nodal features and the graph structure are evaluated, along with their standard deviation.
  • Nodal Features: To evaluate the performance of the reconstructed node features, the Root Normalised Mean Squared Error(RNMSE) is used
@@
    In order to evaluate the performance of reconstructing $\mathbf{X}$ in a graph classification setting, we report the mean RNMSE over all the nodal features in $\mathbf{X}$.
  • Graph Structure: To evaluate the reconstruction performance of the binary adjacency matrix, we use the following metrics
@@
Algorithms 1, 2 and 3 in Appendix A solve an optimization problem using gradient descent. The gradient descent step can be replaced with off-the-shelf optimizers such as Adam or L-BFGS. In all the experiments, Adam is used as the optimizer. In an attack setting, allowing multiple restarts to the optimizer (especially L-BFGS) from different starting points can significantly increase the quality of reconstruction. For a fair comparison, multiple restarts are not used, and all results presented are obtained with a single run of the optimizer. However, in an actual attack setting, an attacker may be able to greatly improve the reconstruction quality by allowing multiple restarts.
In the experiments, a 2-layer GNN model with hidden dimension 100100100 and a sigmoid activation function is employed. The weights of the model are randomly initialized. For all the experiments, the dummy nodal features are initialized randomly, with each entry sampled from the standard normal distribution i.e., $\mathcal{N}(0,1)$. The dummy adjacency matrix is initialized by randomly setting its entries to 0 or 111. The hyperparameters values for the feature smoothness and the sparsity regularizers are set to $\alpha=10^{-9}$ and $\beta=10^{-7}$. These are the values selected by grid search for the choice that leads to the best performance.
@@
We adopt the DLG attack as our baseline and evaluate it across all three settings. A comprehensive comparison with GRAIN is provided in Appendix H. We also extend the experiments involving Node Attacker 1 and Graph Attacker-N to the minibatch setting; these results are detailed in Appendix F.2. As the attack involves numerous hyperparameters, we further analyze a subset of these hyperparameters under various settings, with results presented in Appendix F.3. All reported results represent averages taken over 202020 independent runs of each attack.
@@
We consider the following three datasets:
  • GitHub (Rozemberczki et al., 2021): Nodes represent developers on GitHub and edges are mutual follower relationships. The nodal features are extracted based on the location, repositories starred, employer and e-mail address. Binary labels indicate the user’s job title, either web developer or machine learning developer. The details regarding the number of nodes and edges in this dataset are presented in Table 10.
  • FacebookPagePage (FB) (Rozemberczki et al., 2019): A page-page graph of verified Facebook sites. Nodes correspond to official Facebook pages, links represent mutual likes between sites. Node features are extracted from the site descriptions. The labels denote the categories of the sites. The details regarding the number of nodes and edges in this dataset are presented in Table 10.
  • OGBN-Arxiv (Hu et al., 2020): A large-scale citation network dataset from the Open Graph Benchmark (OGB), representing academic papers from the arXiv repository. Each node corresponds to a paper, and directed edges represent citation links between them. Node features are based on the paper’s title and abstract. The task is node classification: predicting the primary subject area of each paper. As the setting of OGBN-Arxiv aligns with the Node Attacker 1, we evaluate only Node Attacker 1 on this dataset.
@@ 6.1 Node Attacker 1
In this section, we evaluated the performance of feature reconstruction of Node Attacker 1 for the node classification task.
Target node feature reconstruction.
In this experiment, the attacker has no prior knowledge of the graph dataset. Consequently, a 2-layer tree is generated as a dummy graph, with the target node as the root node and each node having degree $d_{tree}=10$. Table 3 reports the RNMSE for reconstructing the target node features, where the attack is performed on 20 randomly selected nodes from each dataset. It can be observed that, for GraphSAGE, the attacker achieves high accuracy in reconstructing the nodal features, whereas for GCN, the attacker struggles to reconstruct the target node features reliably. These observations align with the theoretical insights presented in Propositions 1 and 2. Furthermore, Table 3 also shows the minimum RNMSE achieved among the 20 nodes. While the mean RNMSE for GCN may be large, the presence of smaller minimum RNMSE values indicates that leakage of the nodal features can still occur in practice.
A closer examination of our proposed method, GLG, reveals that it generally attains low RNMSE values for GraphSAGE across the Synthetic, FB, OGBN-Arxiv, and GitHub datasets. For example, on Synthetic and FB, GLG reports mean RNMSE values on the order of $10^{-3}$, indicating a near-exact reconstruction of the target node features. Although the mean RNMSE for GitHub is slightly higher, GLG remains highly effective at recovering node features when GraphSAGE is used. Under the GCN framework, while GLG yields larger RNMSE values overall, its minimum RNMSE can still be as low as 0.180.180.18 on GitHub. This finding underscores that, even in scenarios where GCN poses a more challenging reconstruction problem, GLG can still reveal critical information about certain nodes.
In comparing GLG and DLG, the results highlight the superiority of GLG in most cases. Under the GraphSAGE framework, GLG exhibits significantly lower mean RNMSE on both Synthetic and FB, thereby demonstrating more accurate feature reconstruction. Although DLG achieves a marginally better result on GitHub, this exception does not overshadow the broader trend of GLG’s robust performance. Moreover, under the GCN framework, GLG consistently outperforms DLG on all tested datasets.
Table 4: Reconstruction of $\mathbf{A}$ for Node Attacker 2–G.
@@
ACC AUC AP ACC AUC AP
GLG (Ours)
GraphSAGE 1.00$\pm$0.00 1.00$\pm$0.00 1.00$\pm$0.00 1.00$\pm$0.00 1.00$\pm$0.00 1.00$\pm$0.00
GCN 0.97$\pm$0.08 0.97$\pm$0.02 0.85$\pm$0.04 0.97$\pm$0.02 0.98$\pm$0.001 0.87$\pm$0.02
DLG
GraphSAGE 0.85$\pm$0.14 0.90$\pm$ 0.09 0.65$\pm$0.28 0.83$\pm$0.16 0.87$\pm$0.16 0.60$\pm$0.27
GCN 0.74$\pm$0.25 0.74$\pm$0.26 0.54$\pm$0.20 0.67$\pm$0.28 0.71$\pm$0.27 0.47$\pm$0.27
Table 5: Reconstruction of $\mathbf{X}$ for Node Attacker 2–N.
Framework Facebook GitHub
GLG (Ours)
@@
DLG
[…]
GCN 0.31$\pm$0.19 0.29$\pm$0.24
Detailed results evaluating the performance of node attacker 1 in reconstructing the nodal features of one-hop neighbors are presented in Appendix F.1.
Table 6: Reconstruction of $\mathbf{X}$ (RNMSE) and $\mathbf{A}$ (AUC and AP) with Node Attacker 2-GN, comparing GLG and DLG.
Facebook GitHub
[…]
GLG (Ours)
GraphSAGE 0.0024$\pm$0.003 0.99$\pm$0.01 0.99$\pm$0.01 0.0024$\pm$0.003 0.99$\pm$0.01 0.99$\pm$0.01
GCN 1.39$\pm$0.49 0.97$\pm$0.03 0.75$\pm$0.10 0.58$\pm$0.10 0.99$\pm$0.04 0.83$\pm$0.04
GCN (w/o reg.) 1.19$\pm$1.08 0.95$\pm$0.01 0.65$\pm$0.12 0.81$\pm$1.41 0.94$\pm$0.09 0.72$\pm$0.20
DLG
GraphSAGE 0.003$\pm$0.0 0.99$\pm$0.06 0.99$\pm$0.02 0.0026$\pm$0.03 0.98$\pm$0.02 0.99$\pm$0.1
GCN 1.01$\pm$0.44 0.68$\pm$0.33 0.52$\pm$0.38 0.88$\pm$0.46 0.87$\pm$0.05 0.58$\pm$0.18